Chapter 1 Problems Saved Help Save & Exit Submit Check my work 15 Approximately 14 million Americans are addicted to drugs and alcohol. The federal government estimates that these addicts cost the U.S. economy $300 billion in medical expenses and lost productivity. Despite the enormous potential market, many biotech companies have shied away from funding research and development (R&D) initiatives to find a cure for drug and alcohol addiction. Your firm Drug Abuse Sciences (DAS)- is a notable exception. It has spent $240 million to date working on a cure, but is now at a crossroads. It can either abandon its program or invest another $55 million today. Unfortunately, the firms opportunity cost of funds is 8 percent and it will take another five years before final approval from the Federal Drug Administration is achieved and the product is actually sold. Expected (year-end) profits from selling the drug are presented in the accompanying table. 10 points eBook Print r 9 $0 $0 $O 0 $12,500,000 $14,000,000 $15,700,000 $18,300,000 $20,400,0 Year 2 Year 3 Year 4 r 5 r 6 r 8 What is the net present value of the project? Instruction: Enter your response rounded to the nearest penny (two decimal places). Use a negative sign () where Should DAS continue with its plan to bring the drug to market, or should it abandon the project? Prev 15 of 16 Next>
Answer #1

Note that the $240 million spent to date is irrelevant and it is a sunk cost that will be lost regardless of the decision.We have to check whether the incremental benefits exceed the incremental costs (the $55 million needed to keep the project alive) or not. By using the NPV method, we can execute and compare the investment projects. With this, all costs and benefits that are associated with a project are identified. The Net Present Value of DAS’s R&D initiative is as follows:

NPV = $12,500,000 / (1 + 0.08)5 + $14,000,000 / (1 + 0.08)6 + $15,700,000 / (1 + 0.08)7 + $18,300,000 / (1 + 0.08)8 + $20,400,000 / (1 + 0.08)9 - $55,000,000 =

8,507,452.53 + 8,822,232.02 + 9,160,928.93 +9,887,081.96 + 10,205,102.6 - 55,000,000 =

46,582,798 - 55,000,000 = -$8,417,202



Hence the result is negative, DAS should not spend the $55 million. It should abandon the project.